OpenStudy (anonymous):

In the following unbalanced equation, how many moles of nitrogen monoxide are needed to react with excess ammonia to produce 32.6 grams of water? NH3 + NO --> N2 + H2O

OpenStudy (frostbite):

Start by balancing the equation.

OpenStudy (frostbite):

Just follow the general guide for stoichiometry. :)

OpenStudy (anonymous):

4 NH3 + 6 NO --> 5 N2 + 6 H2O

OpenStudy (frostbite):

Now use the relation for the number of moles you have by a given mass.

OpenStudy (anonymous):

32.6 g H2O/ 18 g H2O = 1.81 mol 1.81 mol H2O * 6 mol NO / 6 mol H2O?

OpenStudy (anonymous):

Is 1.81 the correct answer?

OpenStudy (frostbite):

It sure is. :)

OpenStudy (frostbite):

By just looking at it

OpenStudy (anonymous):

Okay, well thank you for your help :)

OpenStudy (frostbite):

No problem at all... you did it all by your self :)

OpenStudy (frostbite):

Which is the best ever!

OpenStudy (anonymous):

Do you think I'd have to take the molar mass of 6 moles of H2O? Or is what I did okay?

OpenStudy (frostbite):

The 18 g/mol is molar mass? just need to get the units right.

OpenStudy (frostbite):

M(H2O)=2*1 g/mol + 16 g / mol = 18 g/mol

OpenStudy (anonymous):

18 grams was just the mass of 1 mole of H2O, do you think I did it right?

OpenStudy (frostbite):

Well... best you use the unit g/mol as I showed in the calculation.. then your units go perfect :)

OpenStudy (frostbite):

\[n(H _{2}O)=n(NO)\] \[n(NO)=n(H _{2}O)=\frac{ m(H _{2}O )}{ M(H _{2}O) }\] \[M(H _{2}O)=2*M(5)+M(O)=2*1+16=18 g/mol\] \[n(NO)=\frac{ 32.6 g }{ 18 g/mol }=1.81 mol\]

OpenStudy (frostbite):

sorry meant: \[M(H _{2}O)=2*M(5)+M(O)=(2*1+16) g/mol=18 g/mol\]

OpenStudy (anonymous):

Thank you so much! I really appreciate this :)

OpenStudy (frostbite):

lol and not M(5) but M(H) xD

OpenStudy (anonymous):

You're right, though! The units cancel perfectly. I forgot to write mine that way.

OpenStudy (frostbite):

But glad you appreciate it... science teachers love when the units along with the calculations are good :)

OpenStudy (anonymous):

Very true! Thank you again :)

OpenStudy (frostbite):

No problem at all.